OpenStudy (anonymous):

How do I convert this Standard Form equation to Graphing Form? y=3x^2+30x+74

OpenStudy (anonymous):

by completing the square

hero (hero):

1. Subtract 74 from both sides 2. Divide both sides by three 3. Add (10/2)^2 or 25 to both sides 4. Re-write the right side as a binomial square 5. Isolate y once again

hero (hero):

a) \(y - 74 = 3x^2 + 30x\) b) \(\dfrac{y - 74}{3} = x^2 + 10x\) c) \(\dfrac{y - 74}{3} + 25 = x^2 + 10x + 25\) d) \(\dfrac{y - 74}{3} + 25 = (x + 5)^2\) e) \(\dfrac{y - 74}{3} = (x + 5)^2 -25\) \(y - 74 = 3(x + 5)^2 - 75\) \(y = 3(x + 5)^2 - 1\)